A jet airplane is in level flight. The mass of the airplane is ?=8850 kg. The airplane travels at a constant speed around a circular path of radius ?=8.67 mi and makes one revolution every ?=0.115 h. Given that the lift force acts perpendicularly upward from the plane defined by the wings, what is the magnitude of the lift force acting on the airplane?

L=____kN

At what angle is the airplane banked?
Answer #1

Centripetal force F1= mw^2R = 8850*(2pi/(0.115*60*60))^2*8.67*1609 = 28436.6 N

Upward force F2= mg = 8850*9.8 = 86730 N

Lift force L = sqrt(86730^2+28436.6^2) = 91272.8 N = 91.3 kN

angle of banking = arctan(28436.6/86730) = 18.15 degree